OpenStudy (anonymous):

The formula r=2√5L can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid at 60 mph?

OpenStudy (anonymous):

180 is what I get?

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

is the L square? or just the 5?

OpenStudy (anonymous):

rooted that is

OpenStudy (anonymous):

60 = 2 √5L L = 30 / √5 = 13.42 I have assumed that its sqroot of 5 only - not sqroot (5L)

OpenStudy (anonymous):

should 60 be divided by (2/sqt(5))?

OpenStudy (anonymous):

i see yeah that's what I get but it does seam too short

OpenStudy (anonymous):

so not one got that right Length is 180 feet

OpenStudy (anonymous):

problem is that the formula is ambiguous what about the sqroot question we asked?

OpenStudy (anonymous):

if its sqrt of 5L then answer is 180

OpenStudy (anonymous):

is that the answer? if L is sqrt then sqrt(L)=60/(2sqrt(5)) and then squaring L=180

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

think that must be right 13.41 seams way too short.

OpenStudy (anonymous):

is 180 the answer the book gave?

OpenStudy (anonymous):

if so then its sqrt(5L)

OpenStudy (anonymous):

agreed